# NVS-Solver: Video Diffusion Model as Zero-Shot Novel View Synthesizer

### v1

@@ Appendix A Visual Comparison Details
In this section, we visually demonstrate the results as shown in Figs. A-1, A-2, and A-3. The visual result demonstrates the superiority of the proposed method.
@@
Figure A-1: Visual comparison of single view-based NVS results by (a) Text2Nerf [59], (b) 3D-aware [54], (c) MotionCtrl [51], (d) our NVS-Solver (Post). The middle view of each scene highlighted with the red rectangle refers to the input view.
@@
Figure A-2: (a) The two input views of each scene highlighted with the red rectangle. Visual results of multiview-based NVS by (b) 3D-aware [54], (c) MotionCtrl [51], (d) Our NVS-Solver (Post).
@@
Figure A-3: Visual comparison on dynamic scene view synthesis of (a) input frames in the corresponding time of generated images, (b) Deformable-Gaussian [57], (c) 4D-Gaussian [52], (d) 3D-aware [54], (e) MotionCtrl [51], (f) Our NVS-Solver (Post).
Appendix B Quantitative Comparison Details
@@
We give the detailed quantitative comparisons of different methods as shown in Tables B-1, B-2 and B-3. The results demonstrate that the proposed methods, NVS-Solver (DGS) and (Post) outperform SOTA methods in most scenes.
@@
Table B-1: Quantitative comparison of different methods on single view synthesis. For all metrics, the lower, the better.
@@
Scene Metric Sparse Gaussian Sparse Nerf Text2Nerf 3D-aware MotionCtrl NVS-Solver (DGS) NVS-Solver (Post)
@@
Table B-2: Quantitative comparison of different methods on sparse view synthesis. For all metrics, the lower, the better.
Scene Metric Sparse Gaussian Sparse Nerf Text2Nerf 3D-aware MotionCtrl NVS-Solver (DGS) NVS-Solver (Post)
@@
Table B-3: Quantitative comparison of different methods on dynamic view synthesis. ’_1/_2’ mean the we synthesize novel views by moving the camera on the right/left circle trajectory. For all metrics, the lower, the better.
Scene Metric Deformable-Gaussian 4D-Gaussian 3D-aware MotionCtrl NVS-Solver (DGS) NVS-Solver (Post)

### v2

@@ Appendix B Visual Comparison Details
In this section, we visually demonstrate the results as shown in Figs. B-1, B-2, and B-3. The visual result demonstrates the superiority of the proposed method.
@@
Figure B-1: Visual comparison of single view-based NVS results by (a) Text2Nerf (Zhang et al., 2024), (b) 3D-aware (Xiang et al., 2023), (c) MotionCtrl (Wang et al., 2024), (d) Ours (Post). The middle view of each scene highlighted with the red rectangle refers to the input view.
@@
Figure B-2: (a) The two input views of each scene highlighted with the red rectangle. Visual results of multiview-based NVS by (b) 3D-aware (Xiang et al., 2023), (c) MotionCtrl (Wang et al., 2024), (d) Ours (Post).
@@
Figure B-3: Visual comparison on dynamic scene view synthesis of (a) input frames in the corresponding time of generated images, (b) Deformable-Gaussian (Yang et al., 2024c), (c) 4D-Gaussian (Wu et al., 2024a), (d) 3D-aware (Xiang et al., 2023), (e) MotionCtrl (Wang et al., 2024), (f) Ours (Post). ’Failed’ refers to the method cannot work on the condition.
Appendix C Quantitative Comparison Details
@@
We give detailed quantitative comparisons of different methods as shown in Tables C-1, C-2 and C-3. The results demonstrate that the proposed methods, Ours (DGS) and (Post) outperform SOTA methods in most scenes.
@@
Table C-1: Quantitative comparison of different methods on single view synthesis. For all metrics, the lower, the better.
@@
Scene Metric Sparse Gaussian Sparse Nerf Text2Nerf 3D-aware MotionCtrl Ours (DGS) Ours (Post)
@@
Table C-2: Quantitative comparison of different methods on sparse view synthesis. For all metrics, the lower, the better.
Scene Metric Sparse Gaussian Sparse Nerf Text2Nerf 3D-aware MotionCtrl Ours (DGS) Ours (Post)
@@
Table C-3: Quantitative comparison of different methods on dynamic view synthesis. ’_1/_2’ means that we synthesize novel views by moving the camera on the right/left circle trajectory. For all metrics, the lower, the better.
Scene Metric Deformable-Gaussian 4D-Gaussian 3D-aware MotionCtrl Ours (DGS) Ours (Post)
@@
Appendix D Visualization of $\widehat{\lambda}(t,\mathbf{p}_{i})$
In Fig. D-4, we visualize the $\widehat{\lambda}(t,\mathbf{p}_{i})$ of certain sequences in different NVS conditions.
Figure D-4: Visualization of $\widehat{\lambda}(t,\mathbf{p}_{i})$ for (a) single view synthesis, (b) sparse view synthesis (c) dynamic view synthesis.
Appendix E Warping Strategy
We illustrated the warping strategies for different NVS conditions in Fig. E-5.
Figure E-5: Illustration of different warping patterns. From left to right, there are single-view, monocular video, and multi-view-based NVS.
Appendix F Performance of Our Method across Different Depth Estimation Methods
We conducted experiments on our method using various depth estimation approaches, including DINOV2 (Oquab et al., 2023) and DepthAnything V2 (Yang et al., 2024b). The results, shown in Table F-4, clearly demonstrate that our method consistently performs well across different depth estimation techniques, surpassing the state-of-the-art NVS method, MotionCtrl (Wang et al., 2024). As expected, better depth maps indeed lead to more accurate NVS, as evidenced by the comparison between DepthAnythingV1 (Yang et al., 2024a) and DINOV2 Oquab et al. (2023). Furthermore, the performance of DepthAnythingV2 is only comparable to DepthAnythingV1 due to one scene that appears to be an outlier. Upon removing this scene (marked with *), DepthAnythingV2 significantly outperforms DepthAnythingV1. Additionally, DINOV2 achieves a lower FID due to the amplification of rendering pose errors, which facilitates easier reconstruction. Fig. F-6 visualizes the results of our method with DepthAnything V1, V2, and DINOV2. These comparisons confirm the robustness of our method with respect to depth estimation techniques, consistently delivering high performance across different depth estimation modules, including DINOV2, DepthAnything V1, and V2.
Table F-4: Quantitative comparison of our method with DepthAnythingv1   DepthAnythingv2, and DinoV2.For all metrics, the lower, the better.
Methods FID ATE RPE-T RPE-R
MotionCtrl (SOTA) 179.24 3.851 0.705 0.835
Ours + DepthAnythingv1 165.12 0.767 0.156 0.170
Ours + DepthAnythingv2 162.896 0.831 0.110 0.170
Ours + DepthAnythingv2* 163.93 0.243 0.056 0.056
Ours + DinoV2 158.24 2.395 0.454 0.674
Figure F-6: Visual results of our method with (a) DepthAnythingv1, (b) DepthAnythingv2,(c) DinoV2.
Appendix G Comparisons with Image Inpainting Results
Fig. G-7 visualizes the comparison between using an image inpainting method and our approach. Here, we use SDEdit (Meng et al., 2022) as the image inpainting method. When per-view inpainting is used to render novel views, it becomes difficult to ensure that the synthesized views are naturally consistent. As expected, the visual comparisons show that consistency across different views cannot be guaranteed when solving NVS as a per-view inpainting task with SDEdit.
Figure G-7: Visual comparisons of a synthesized video sequence by (a) image inpainting method, and (b) our method.
Appendix H 360° NVS Strategy
For a single view input, we first apply the proposed method to rotate 120° to the left and right sides. Moreover, to achieve consistency from different directions, we further process the remaining 120° in a multi-view prompt manner, i.e., we warp both side views to the central as prompt images. Note that in the first two 120° rendering processes, it’s also difficult to directly achieve such a large range NVS. Thus, we first reconstruct 30° NVS with 24 frames. Then, we sample 12 frames from the reconstructions as the first 12 prompt images for 60° NVS. Next, we can achieve 24 frames of 120° NVS by sampling 12 prompts from 60° NVS reconstruction. When given multiple views, the proposed method achieves 360° NVS by treating two neighboring images as side views of a pair, warping them towards the central as prompt images.
Appendix I Quantitative Comparison on LPIPS
We give detailed quantitative comparisons on LPIPS of different methods as shown in Tables I-5. The results demonstrate that the proposed methods, Ours (DGS) and (Post) outperform SOTA methods in most scenes.
Table I-5: Quantitative comparison on LPIPS of different methods on single view synthesis. For all metrics, the lower, the better.
Scene Metric Sparse Gaussian Sparse Nerf Text2Nerf 3D-aware MotionCtrl Ours (DGS) Ours (Post)
Auditorium LPIPS – – – – 0.622 0.707 0.598 0.563 0.567
Barn LPIPS – – – – 0.428 0.643 0.443 0.427 0.438
Church LPIPS – – – – 0.642 0.714 0.627 0.623 0.634
Family LPIPS – – – – 0.647 0.794 0.540 0.523 0.525
Ignatius LPIPS – – – – 0.634 0.853 0.586 0.542 0.536
Palace LPIPS – – – – 0.765 0.627 0.577 0.554 0.549
Appendix J Mesh reconstruction
We have applied 2D Gaussian Splatting to reconstruct the mesh on the generated 360° scene. As demonstrated in Fig.J-8, our method successfully maintains geometric consistency across the generated images.
Figure J-8: Mesh reconstruction of synthesized 360° NVS.
Appendix K Comparison result on Dycheck
We conducted comparison experiments on three scenes from the Dycheck dataset. Table. K-6 illustrates the results of our method alongside two Gaussian-based methods. The three scenes from Dycheck were captured using two cameras, each positioned at a considerable distance from the other. In our experiments, we used a monocular video from one camera as input and generated a video following the trajectory of the other camera. This setup introduces a significant content unoverlap challenge, as the input and target videos capture substantially different perspectives of the scene, making the task particularly difficult. To address potential scale inconsistencies, we utilized the depth maps provided in the dataset. Our method outperformed the 4D-Gaussian approach across all three metrics.The Deformable-Gaussian method failed to produce viable results in this challenging scenario.
Table K-6: Quantitative comparison of our method with 4D-Gaussian, Deformable-Gaussian on Dycheck. $\uparrow$ (resp. $\downarrow$) means the larger (resp. smaller), the better.
Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Deformable-Gaussian
4D-Gaussian 12.68 0.346 0.737
Ours 15.84 0.385 0.410
Appendix L Ablation on different weight function
We conducted ablation experiments to evaluate different weight function choices, including (1) a constant value of 0.5
$\widehat{\lambda}(t,\mathbf{p}_{i})=0.5,$ (36)
(2) a linear function
$\widehat{\lambda}(t,\mathbf{p}_{i})=t,$ (37)
(3) an exponential function
$\widehat{\lambda}(t,\mathbf{p}_{i})=e^{(t+1)}.$ (38)
The results are presented in Table. L-7 and Figure. L-9. These results highlight the superiority of our weight design, as all three simple weight functions lead to decreased performance in generated image quality and trajectory accuracy, particularly when using a constant value for weighting.
Table L-7: Quantitative comparison of different weight functions. For all metrics, the lower, the better.
Methods FID ATE RPE-T RPE-R
Constant (0.5) 175.68 6.09 1.06 0.864
Linear 174.23 1.04 0.210 0.199
Exponential 174.60 1.363 0.312 0.330
Ours 165.12 0.767 0.156 0.170
Figure L-9: Visual comparison of single view-based NVS results utilizing weight function as (a) constant, (b) Linear, (c) Exponential, (d) Ours (Post).
Appendix M Comparison with ZeroNVS
We present a comparison of 360-degree NVS comparison results in Figure M-10. The results highlight significant differences between the approaches. ZeroNVS generates videos by adhering to a specific and relatively constrained pattern, which limits the diversity and complexity of the generated scenes. In contrast, our method demonstrates the ability to produce videos with more intricate and realistic background geometry.
Figure M-10: Comparison on 360° NVS results between (a) Ours and (b) ZeroNVS (Sargent et al., 2024).
Appendix N Comparison with Photoconsistent-NVS
We conducted a comparative experiment on Photoconsistent-NVS, and the quantitative results are presented in Figure N-11.
Figure N-11: Visual comparison between (a) Ours and (b) Photoconsistent-NVS (Yu et al., 2023).
Appendix O Comparison with Sparse Gaussian on Extreme Zoom In and Out
We conducted a comparative experiment by performing extreme zoom-in and zoom-out operations on Sparse Gaussian (Xiong et al., 2023), as illustrated in Figure O-12.
Figure O-12: Visual comparison between (a) Ours and (b) Sparse Gaussian (Xiong et al., 2023), where for the first two scenes both methods are inputted with the same one view. Moreover, we also input the 12 views to Sparse Gaussian on the third scene and keep ours with 1 scene, which is “unfair” to our method. Experimental results demonstrate the robustness and consistency of our method with large camera pose changes.